### v1

# InquireMobile: Teaching VLM-based Mobile Agent to Request Human Assistance via Reinforcement Fine-Tuning

Abstract
Recent advances in Vision-Language Models (VLMs) have enabled mobile agents to perceive and interact with real-world mobile environments based on human instructions. However, the current fully autonomous paradigm poses potential safety risks when model understanding or reasoning capabilities are insufficient. To address this challenge, we first introduce InquireBench, a comprehensive benchmark specifically designed to evaluate mobile agents’ capabilities in safe interaction and proactive inquiry with users, encompassing 5 categories and 22 sub-categories, where most existing VLM-based agents demonstrate near-zero performance. In this paper, we aim to develop an interactive system that actively seeks human confirmation at critical decision points. To achieve this, we propose InquireMobile, a novel model inspired by reinforcement learning, featuring a two-stage training strategy and an interactive pre-action reasoning mechanism. Finally, our model achieves an 46.8% improvement in inquiry success rate and the best overall success rate among existing baselines on InquireBench. We will open-source all datasets, models, and evaluation codes to facilitate development in both academia and industry.
{strip}[Uncaptioned image]
Figure 1: An example of a high-stakes scenario involving irreversible file deletion, which requires human confirmation before execution. In fact, situations requiring human assistance are widespread.
1 Introduction
Recent advances in Vision-Language Models (VLMs) have significantly propelled the capabilities of mobile agents, enabling them to autonomously perceive and interact with complicated real-world environments based on human instructions (Hurst et al. 2024; Bai et al. 2025; Li and Huang 2025). Traditionally, these mobile agents (Hong et al. 2024; Qin et al. 2025) adopt a fully autonomous paradigm: after receiving user instructions, the mobile agent independently interacts with its mobile environment to accomplish designated tasks.
However, this paradigm presumes “absolute trust” in the agent’s decision-making abilities—a presupposition that becomes problematic when model understanding or reasoning capacity is insufficient. In such cases, unverified autonomy can result in severe consequences, especially in high-stakes scenarios such as online payments, order placements, or the handling of sensitive personal data as shown in Figure 1. For instance, a misinterpreted instruction at a payment interface could lead to unintended financial transactions.
Recognizing these limitations, we argue that robust mobile agents should not only interact with their environment but must also establish a proactive feedback mechanism with users. Specifically, agents should be equipped to seek human clarification or confirmation at critical junctures, especially when the context suggests potential risk or when the agent’s confidence is low. This approach effectively cuts the blind reliance on model outputs and incorporates a human-in-the-loop safeguard, enhancing the overall safety, transparency, and trustworthiness of mobile agent systems.
Name Eval Mode # Tasks # Apps Language Inquire Data Online
AITW (Rawles et al. 2023) static - - EN
AndroidControl (Li et al. 2024) static - - EN
AMEX (Chai et al. 2024) static - - EN
A3 (Chai et al. 2025) dynamic 201 21 EN
AppAgent (Zhang et al. 2025a) dynamic 45 9 EN
AndroidArena (Xing et al. 2024) dynamic 221 14 EN
AndroidWorld (Rawles et al. 2024) dynamic 116 20 EN
Android-Lab (Xu et al. 2024) dynamic 138 9 EN
Mobile-Env (Zhang et al. 2023) dynamic 224 15 EN
CAGUI (Zhang et al. 2025b) static 603 - CN
InquireBench (Ours) dynamic 173 37 EN&CN
Table 1: Comparison of several existing GUI agent benchmark. Among them, “Eval Mode” denotes the evaluation protocol: static (single screenshot with ground-truth history) or dynamic (step-by-step interaction). “Online” denotes whether the app requires a network connection, e.g., offline built-in system apps or online commercial apps.
To bridge this research gap, we introduce InquireBench, a novel benchmark specifically designed to evaluate a mobile agent’s ability to inquire and interact safely with users. Specifically, we systematically construct diverse and challenging scenarios by leveraging both existing instructions and a vast new set of instructions generated by GPT-4o. By simulating “random walks” on real mobile phones, we trigger a breadth of potential inquire cases across 5 categories (i.e., intent confirmation, privacy and security, risk scenarios, combination and others), for evaluating user-agent interactive performance. Surprisingly, comprehensive evaluation of both open-source and closed-source VLM-based agents (Bai et al. 2025; Luo et al. 2025; Lu et al. 2025a) on InquireBench reveals a near-zero baseline performance, underscoring the urgent need for new techniques tailored for user-agent collaborative situations.
To achieve this goal, we propose InquireMobile, a model designed to teach VLM-based mobile agents to request human assistance through reinforcement fine-tuning. This new mobile agent model employs a two-stage training strategy: it begins with supervised fine-tuning (SFT) for robust format acquisition, and is followed by Group Relative Policy Optimization (GRPO) (Shao et al. 2024) training to enhance the model’s reasoning and thinking capabilities. Crucially, InquireMobile incorporates an interactive pre-action reasoning mechanism, where the agent proactively inquires from the user before executing critical actions—enabling informed decision-making through human-in-the-loop interaction. In our experiments, InquireMobile achieves a significant performance gain of 46.8% points in inquiry success rate and the best success rate over existing baselines on InquireBench, demonstrating the feasibility and necessity of proactive user engagement in agent-driven automation.
The contributions are summarized as follows:
  • 1) We systematically address a crucial yet underexplored situation—the agent’s capacity to inquire and collaborate with users in mobile environments;
  • 2) We introduce InquireBench, a benchmark designed to evaluate the ability of mobile agents to effectively interact with users when human intervention is required.
  • 3) We present InquireMobile, a model that achieves state-of-the-art results via a two-stage training strategy.
2 Related Work
2.1 VLM-based GUI Agents
With advances in computer vision and vision-language models (VLMs), agents can directly utilize screen-visual information—such as screenshots—to better perceive and interact with the on-screen environment. Recent studies have shifted from complex framework designs to exploring end-to-end agent strategies based on VLMs, enhancing the perception and planning capabilities of VLMs. For example, CogAgent (Hong et al. 2024) presents a vision-language foundation model specializing in GUI understanding and planning, relying solely on visual inputs. UI-TARS (Qin et al. 2025) continually enhances Qwen-2-VL (Wang et al. 2024) in perception, reasoning, and memory by training on approximately 50 billion tokens. OS-Kairos (Cheng et al. 2025) employs a confidence-driven strategy to seek help from advanced models or humans when actions are unreliable, ensuring smooth task progress.
Figure 2: Data Collection Pipeline of our InquireBench. Among them, we employ a random walk approach to trigger the potential inquiry scenario, in which the agent seek human assistance.
2.2 RFT-tuned GUI Agents
Inspired by the reinforcement fine-tuning (RFT) approach of DeepSeek-R1 (Guo et al. 2025), RFT-tuned VLMs (Zhou et al. 2025a; Liu et al. 2025b; Chen et al. 2025) have demonstrated excellent performance across various vision tasks. By combining SFT with reinforcement learning (e.g., GRPO), rule-based reward mechanisms significantly enhance the models’ reasoning and generalization capabilities. UI-R1 (Lu et al. 2025b) and GUI-R1 (Luo et al. 2025) extend this strategy to GUI agents, employing simple rule-based reward functions to evaluate the correctness of actions, achieving notable performance improvements with only limited datasets. InfiGUI-R1 (Liu et al. 2025a) proposes a reasoning-oriented, two-stage training paradigm, designed to gradually evolve the agent from a reactive actor to a deliberate planner. GUI-G1 (Zhou et al. 2025b) enhances the GUI interaction ability of VLMs through unified action-space rule modeling. Mobile-R1 (Gu et al. 2025) further introduces an interactive multi-round reinforcement learning framework with task-level rewards, significantly improving the agent’s exploration and error-correction abilities.
Although these models have achieved impressive performance on existing benchmarks (Cheng et al. 2024; Li et al. 2024, 2025), they focus on action execution while neglecting the interactive feedback from users. This poses significant challenges to the user privacy and security of GUI agents.
3 InquireBench
We introduce InquireBench, a novel benchmark specifically designed to evaluate a mobile agent’s ability to inquire and interact safely with users.
Specifically, we first collect 80,345 raw data samples by simulating “random walks” on real mobile phones. A portion of this data is used to construct the benchmark inquiry GUI data, while the remaining samples are used for general GUI training data construction. Finally, InquireBench comprises 975 annotated data collected from 173 tasks. The pipeline of data collection is shown in Figure 2, which is divided into situation triggering (Section 3.1.1) and thinking generation (Section 3.1.2).
3.1 Inquiry GUI Data
3.1.1 Triggering the Inquiry Situation
Existing datasets are typically constructed under idealized conditions and fail to capture the diverse range of anomalies encountered in real-world deployment. Traditional dataset construction methods typically rely on manual processes, including task design, execution planning, and human demonstration. However, such approaches are not suitable for our benchmark, as the interactive steps in our setting involve inherent randomness and unknown factors. Inspired by RevAct (Yang et al. 2025), we adopt a novel data collection strategy that reverses the conventional workflow and enables semi-automated dataset generation. Specifically, we first collect a large number of screenshots by simulating “random walks” on real mobile phones [^1] The devices used for data collection include Huawei and Xiaomi phones, both equipped with Android systems., which better reflects the natural interactions observed in daily life. Notably, to account for the significant differences in categories and usage patterns between Chinese and English apps, users are instructed to interact with each type of app separately. Each app is then evaluated within its respective system language setting.
After the “random walks”, we obtain approximately 80,345 screenshots. Human annotators are then asked to identify the interactive screenshots that require human intervention when agents encounter these screens, such as account login or payment execution. The interactive scenarioed are summarized into five categories:
  • Intent Confirmation: When the agent is unable to determine the next action (e.g., pop-up windows, advertisements, or ambiguous user instructions), it proactively seeks guidance from the user.
  • Privacy and Security: Actions such as login or permission granting require explicit user confirmation.
  • Risk Scenarios: Cases involving advertisement pop-ups, payment operations, or file deletions that necessitate user intervention.
  • Combination: Involving multiple types of interaction reasons or scenarios.
  • Others: Scenarios that cannot be classified into the above four categories.
Human annotators are instructed to determine whether each screenshot requires user interaction and to assign it to the corresponding interactive category. Screenshots that do not require user intervention are discarded. Because interactive scenarios are difficult to trigger, we ultimately collect a total of 975 images that require user actions.
3.1.2 Thinking Generation
After obtaining the images labeled with interactive categories, we proceed to generate appropriate user tasks and corresponding thinking, i.e., instructions and responses, for each image.
Valid Task
Given an image and its labeled interactive category, the annotators are asked to write a valid task for the image. Valid tasks are strictly defined according to two criteria: 1) Executable on the Device: The user task must be executable on a real device, meaning that the task cannot require actions beyond the app’s actual capabilities. For example, purchasing items via Spotify or subscribing to a non-existent membership type in Tencent Video are not allowed. 2) Consistent with the Interaction Type: The user task must correspond to the interactive category assigned to the image.
Figure 3: Distribution of our InquireBench dataset. The top three most frequent apps are listed for each category.
Interactive Thinking
Based on each image and its corresponding valid task, we prompt GPT-4o-0806 to generate both the interactive reason and the interactive content for each image-instruction pair. The interactive reason describes the agent’s internal reasoning in the current situation, explaining why user intervention is required. We then require human annotators to review and, if necessary, revise the thinking content. The annotation procedure is as follows: annotators first read the instruction and the corresponding screenshot, then review the generated interactive reason and interactive content.
Annotators assess whether each item is appropriate and reasonable for the given context, and then decide whether to retain the original content, edit any unreasonable parts, or rewrite them entirely if necessary. They evaluate the model’s “thinking” for reasonableness using the format shown in Figure 4. Redundant or incorrect data is rewritten.
Figure 4: Interactive thinking format. Red denotes the user instruction, blue indicates the current state or interface, and purple shows the reason and category in inquiry scenarios.
Inquiry Content
The interactive content refers to the external message presented to the user, specifying the exact content used to interact with them. Annotators will revise it to eliminate redundancies and ensure that it is not offensive.
3.1.3 Statistics
The category distribution is shown in Figure 3. More details can be found in Appendix 7.3.
3.2 General GUI Data
Moreover, to enhance the model’s inquiry ability while retaining its general GUI capability, we additionally collected 3,000 general-GUI data that do not involve interactive requirements from the original 80,345 raw data samples. We then selected complete trajectories and asked human annotators to first write an appropriate instruction for each trajectory, and subsequently annotate every step with the corresponding golden thoughts and actions.
4 InquireMobile
Figure 5: Our training framework consists of two stages: an initial cold start stage with supervised fine-tuning, followed by online GRPO training using rule-based rewards.
Our InquireMobile adopts a two-stage training strategy: (1) format finetuning by supervised fine-tuning (SFT), and (2) inquiry enhancement via GRPO training with verifiable rewards, to improve the agent’s interactive capabilities, as shown in Figure 5. The data used in the two training stages include both inquiry-related and general GUI human-annotated data, as described in Section 3. Given a task presented with a user instruction in natural language, the mobile agent is responsible for making action decisions to complete this instruction on the phone, if necessary, calls the user for asistance or clarification.
4.1 Preliminary
Rule-based RL can enhance the reasoning capabilities of multimodal large language models (MLLMs) through policy-based algorithms such as Group Relative Policy Optimization (GRPO) (Shao et al. 2024). By using group-normalized, token-level advantages, GRPO lowers reward sparsity and variance while removing the need for a separate value function or critic network.
In GRPO, the model generates a set of $N$ candidate responses $O=\left\{o_{1},o_{2},...,o_{N}\right\}$ for each task. Each response is evaluated by taking the corresponding actions and computing its reward $\left\{r_{1},r_{2},...,r_{N}\right\}$. Unlike PPO (Schulman et al. 2017), which relies on a single reward signal and a critic to estimate the value function, GRPO normalizes these rewards to calculate the relative advantage of each response. The relative advantage $A_{i}$ of the $i$-th response is calculated as follows:
$A_{i}=\frac{r_{i}-\mathbf{mean}(\left\{r_{1},r_{2},...,r_{N}\right\})}{\mathbf{std}(\left\{r_{1},r_{2},...,r_{N}\right\})}$ (1)
where $\mathbf{mean}$ and $\mathbf{std}$ denote the mean and standard deviation of the rewards, respectively. Given a batch of $G$ generated responses $\{o_{i}\}_{i=1}^{G}$ from a task, the GRPO objective function is defined as:
$\begin{split}J_{\text{GRPO}}(\theta)&=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\min\left[\frac{\pi_{\theta}(o_{i}(t)|o_{i},<t)}{\pi_{\text{old}}(o_{i}(t)|o_{i},<t)}\hat{A}_{i,t},\right.\\ &\quad\left.\text{clip}\left(\frac{\pi_{\theta}(o_{i}(t)|o_{i},<t)}{\pi_{\text{old}}(o_{i}(t)|o_{i},<t)},1-\epsilon,1+\epsilon\right)\hat{A}_{i,t}\right]\end{split}$ (2)
where $\pi_{\theta}$ and $\pi_{old}$ are the current and old policy. $\epsilon$ is the clipping hyperparameter. $\hat{A}_{i,t}$ is the group-normalized advantage for token $o_{i}(t)$ in response $o_{i}$.
4.2 Stage 1: Supervised Fine-tuning
The model was first trained using Supervised Fine-tuning (SFT) to endow it with the ability to produce structured outputs and to acquire fundamental GUI interaction skills. The training data includes both inquiry data and general GUI data, as described in Section 3.1 and Section 3.2.
4.3 Stage 2: Rule-Based RL Training
Subsequently, the model was trained using GRPO with verifiable rewards. The reward function consists of format reward and action reward, defined as:
$R=R_{F}+R_{T}+R_{A},$ (3)
where $R_{A}$ and $R_{T}$ quantify the correctness of the action arguments and action type, respectively, thereby ensuring the overall correctness of the executed action. $R_{F}$ ensures that the output adheres to the expected structural format.
Format Reward ($R_{F}$)
Following previous work (Meng et al. 2025; Gu et al. 2025; Huang et al. 2025), we introduce the format reward $R_{F}$ to encourage the model to generate structured and interpretable outputs.
  • <think>: The internal reasoning process.
  • <tool_call>: The final answer in the form of a JSON object containing the function name and arguments.
Moreover, $R_{F}$ is set to $1$ to encourage format matching and to 1-1 for stricter penalties on errors.
Action Type Reward ($R_{T}$)
$R_{T}$ assesses the the correctness of the predicted action type. In our tasks, the action space consists of nine types, including general actions such as click and swipe, as well as the inquiry action call_user. The action space is detailed in Table 5 in Appendix 7.2. $R_{T}$ is 1 if the action type exactly matches the ground truth otherwise, it is 0.
Action Argument Reward ($R_{A}$)
$R_{A}$ assesses the correctness of the predicted action argument. The reward is computed differently depending on the action type.
  • For coordinate-based actions (e.g., click, swipe), $R_{A}$ is $1$ if the predicted coordinate $C=[x,y]$ falls within the ground truth bounding box $B=[x_{1},y_{1},x_{2},y_{2}]$ of the target GUI element; otherwise, it is 0, to ensure precise interaction.
  • For text-based actions (e.g., type, call_user), we compute the BLEU (Papineni et al. 2002) score between the generated text and the ground truth. The BLEU score is normalized to the range $[0,1]$.
Method Chinese English Average
ISR SR Score ISR SR Score ISR SR Score
AppAgent Framework
Qwen2.5-VL-72B - 6.7% - - 4.2% - - 5.45% -
GPT-4o-0806 - 2.1% - - 4.2% - - 3.15% -
Gemini-2.5-pro - 1.0% - - 6.3% - - 3.65% -
Claude-3.5-Sonnet v2 - 8.4% - - 4.2% - - 6.30% -
Mobile-Agent-E Framework
Qwen2.5-VL-72B - 1.0% - - - - 0 0.50% -
GPT-4o-0806 - 2.1% - - 1.0% - - 1.55% -
Gemini-2.5-pro - 4.2% - - 4.2% - - 4.20% -
Claude-3.5-Sonnet v2 - 3.2% - - 4.2% - 0 3.70% -
Model-based Framework
UI-R1-3B 3.2% 5.3% 0.54 6.3% 6.3% 0.73 4.75% 5.80% 0.64
UI-R1-3B-E 3.2% 4.2% 0.41 2.1% 3.2% 0.54 2.65% 3.70% 0.48
GUI-R1-3B 10.5% 9.5% 0.91 1.1% 0 0.37 5.80% 4.75% 0.64
Qwen2.5-VL-3B 2.1% 4.2% 0.33 3.2% 1.1% 0.42 2.65% 2.65% 0.38
+ InquireMobile Stage1 17.9% 3.2% 0.24 37.9% 1.1% 0.27 27.9% 2.15% 0.26
+ InquireMobile Stage2 5.3% 5.3% 0.69 5.3% 6.3% 0.65 5.30% 5.80% 0.67
+ InquireMobile Stage1 & Stage2 49.5% 10.5% 0.83 55.8% 5.3% 0.72 52.6% 7.90% 0.78
Table 2: Main results on InquireBench. ISR denotes the inquiry success rate and SR denotes the task success rate. Bold and underline indicate the best and second-best results, dash (“–”) indicates that a result is not available.
5 Experiment
5.1 Implementation Details
Testing Environment
The Android Studio emulator and two physical Android mobile phones serve as our primary mobile GUI interaction environments. Local monitoring scripts run on the host machine and connect to each device to manage the interaction loop.
Datasets and Benchmark
In the two training stages, we utilized 975 inquiry data to enhance interactive ability and 3,000 general GUI data to retain model’s mobile agent ability. Further more, we constructed a bilingual benchmark of 190 carefully curated instructions—95 Chinese and 95 English—each crafted to elicit uncertainty and require human-in-the-loop interaction. To mirror authentic mobile usage, each task is paired with a realistic execution environment, as detailed in Appendix 7.4.
Training Settings
Our experiments utilize Qwen2.5-VL-3B as the base model, with GRPO implementation. The training procedure consists of two stages trained on 4 H100 GPUs. In Stage 1, we perform supervised fine-tuning (SFT) with LoRA (lora_rank =8=8) for 2 epochs with a learning rate of $1.0\times 10^{-4}$. In Stage 2, GRPO is applied for 2 epochs with 4 generations per sample and a temperature of 1 for exploration. The max steps for each task is set to 15.
Baselines
AppAgent (Zhang et al. 2025a), Mobile-Agent-E (Wang et al. 2025), Qwen2.5-VL-3B (Bai et al. 2025), GUI-R1-3B (Luo et al. 2025), UI-R1-3B and UI-R1-3B-E (Lu et al. 2025a) are baselines.
Evaluation Metrics
To evaluate the performance of the existing frameworks and models in real-world scenarios, we use a Huawei phone instead of simulators as the testing environment. We evaluate the agent’s performance on these tasks using the following metrics:
  • Task Success Rate (SR): the proportion of tasks the agent completes successfully.
  • Inquiry Success Rate (ISR): the proportion of cases in which the agent makes a correct inquiry to the user at the appropriate time.
  • Task Completion Score (Score): the score of task completion rated by GPT-4o [^2] The version we used is GPT-4o-0806. Prompts provided to GPT-4o can be found in the Appendix 7.1..
5.2 Experimental Result
Figure 6: Comparison of reasoning trajectories between InquireMobile and Qwen2.5-VL-3B-Instruct on the task “One of my socks is torn”. In this case, Qwen2.5-VL-3B failed at the first step, while InquireMobile completed the whole task accurately.
As shown in Table 2, we have the following observations:
1) Our model outperformed all baselines on average across all metrics, especially achieving an ISR of 52.6%, which is 46.8 points higher than the best baseline. With two-stage training, the model successfully mastered the ability to request human assistance in interactive scenarios.
2) We observed that both the success rate (SR) and the trajectory score of all baselines were very low. Specifically, after undergoing training in stages 1 and 2, InquireMobile achieved a success rate of 7.9% and a trajectory score of 0.78. Despite our InquireMobile achieving state-of-the-art performance compared to existing baselines, its overall performance in task completion and trajectory evaluation is still unsatisfactory. We attribute these results to two main factors: i) the limited inquiry ability of the models, ii) the complex and realistic scenario settings, as detailed in Appendix 7.4. It highlights a gap between real user instructions and dynamic environments.
  • Grounding & navigation. Agents frequently issue invalid click actions, revealing insufficient grounding in real mobile contexts and within dynamic online applications.
  • App-level priors. Agents sometimes open the wrong app first because they lack basic knowledge, real-world of commercial app identities.
  • In-app page comprehension. Even in the right app pages, agents struggle to understand the page or icons, causing errors in task progress and subsequent erroneous actions.
Figure 7: Reward during Stage 2 training.
3) After Stage 1 training (SFT-only), the model has a higher Inquiry-Success Rate (ISR) than Stage-2 alone, but its overall Success Rate (SR) and trajectory scores are much lower. Manual review shows the agent in Stage 1 often asks the user for confirmation unnecessarily, even when the user wants it to proceed independently. These unnecessary queries introduce redundant actions and lead to low-quality exploration paths. The root cause is that SFT stage doesn’t expose the policy to real-device complexities or distinguish between necessary and unnecessary inquiries.
4) Our two-stage training approach, InquiryMobile, strikes a better balance. It retains the general GUI skills learned in Stage 2, while introducing inquiry capabilities in Stage 1—specifically when ambiguity, risk, or a critical need for intent clarification arises. this approach ensures inquiries occur only when necessary. Therefore, the two stages complement each other: Stage 1 guides interactions, while Stage 2 curbs unnecessary inquiries, leading to higher task success and better user interactions.
Moreover, the reward score of Stage 2 training in Figure 7 exhibits a slow upward trend with fluctuations, suggesting gradual learning despite some instability.
5.3 Human Satisfaction
To further investigate whether this “inquire” strategy meets human expectations, we recruited approximately 200 participants from various public places, including stations, schools, and shopping malls. These participants rated our system on a five-point scale ranging from 0 to 5, with 5 indicating the highest level of satisfaction.
The results are reported in Table 3. Our InquireMobile (Stage 1 & Stage 2) achieves the best performance in both the English and Chinese settings and is the only model that attains a satisfaction score above 2. After the two-stage training, InquireMobile can dynamically request human assistance based on the user’s instructions and the current state. Although GUI-R1-3B performs well in task completion and reaches an average satisfaction score of 1.8, it shows weak inquiry capability. By contrast, InquireMobile Stage 1 issues many inquiry actions; however, participants noted that its repeated call_user requests were redundant and did not contribute to task completion, which is consistent with the findings in Section 5.2.
Method Human Satisfaction
EN CN Avg
UI-R1-3B 1.7/5 1.8/5 1.75/5
UI-R1-3B-E 1.2/5 1.4/5 1.30/5
GUI-R1-3B 1.9/5 1.7/5 1.80/5
Qwen2.5-VL-3B 1.0/5 1.2/5 1.10/5
+ InquireMobile Stage1 1.4/5 1.5/5 1.45/5
+ InquireMobile Stage2 1.6/5 1.3/5 1.45/5
+ InquireMobile Stage1 & Stage2 2.3/5 2.2/5 2.25/5
Table 3: Satisfaction results on InquireMobile, where five-point scale ranging from 0 to 5.
5.4 Qualitative Visualization
We randomly selected several examples from the test set for qualitative analysis. As illustrated in Figure 6, we have made the following observations: 1) Due to ambiguous user instructions, Qwen2.5-VL-3B struggled to comprehend the task intent and became stuck at the initial step. 2) In contrast, InquireMobile demonstrated robust interactive reasoning capabilities when faced with unclear instructions. Instead of halting, it proactively engaged the user to clarify the task objectives, accurately inferred the underlying intent, and efficiently completed the task in a concise manner.
6 Conclusion
In this paper, we present InquireBench and InquireMobile, establishing a new paradigm for agent-human-environment interaction in mobile scenarios. Our extensive experiments reveal that existing VLM-based agents achieve near-zero performance when faced with scenarios requiring human interaction, while our InquireMobile successfully addresses these limitations with an 46.8% improvement in inquiry success rate, the best overall success rate and trajectory score. This significant performance gain validates our core thesis that proactive user engagement is crucial for safe and effective mobile agent deployment.